# Exploring Graph Representations of Logical Forms for Language Modeling

### v1

@@ 5 Downstream Tasks
In order to demonstrate the viability of LFLMs, we evaluated GFoLDS on a series of downstream tasks: our primary objective was to demonstrate that language models over logical forms are able to learn from less data than their textual counterparts. To that end, we compared GFoLDS to the BERT comparison (BERT-C) models introduced in Section 4.1, in order to show that GFoLDS is able to outperform a textual LM when pretrained on the same amount of data: it then follows that GFoLDS would be able to perform as well as a textual LM with less pretraining data.
We also compared GFoLDS to the original $\text{BERT}_{\text{base}}$ and $\text{BERT}_{\text{large}}$ (uncased) models, both of which are pretrained on roughly 6.5 times more data than GFoLDS.
@@ 5.1 Datasets
We evaluated the models on four downstream tasks: the RELPRON task described in Section 4, SNLI Bowman et al. (2015), the MegaVeridicality V2.1 factuality dataset White et al. (2018), and the McRae et al. (2005) property inference dataset.
@@ 5.1.1 SNLI
All five models were fine-tuned end-to-end on SNLI with a two-layer, feed-forward MLP classifier over their mean-pooled token/node embeddings. The details of our SNLI fine-tuning hyperparameters are located in Appendix E.
Graph Representations
@@
$\text{BERT}_{\text{large}}$ 0.768 0.769 91.1% 85.6% 0.241
$\text{BERT}_{\text{base}}$ 0.667 0.690 90.7% 84.2% 0.247
@@
Table 1: Results for GFoLDS and the four BERT models on downstream tasks.
@@
The objective in this task is to determine whether the event denoted by a subordinate clause is true, given the context of the matrix clause (see Table 3 in the Appendix). The MegaVeridicality V2.1 dataset consists of 5,026 examples, each with ten annotations of the factuality of the event: for this experiment, we converted the data into a binary classification task (see Appendix F).
@@
To evaluate the quality of GFoLDS’ embeddings in comparison to BERT, we evaluated the models on a property inference task using the McRae et al. (2005) feature norm database. This database consists of a set of 541 concepts (words) $W$ and 2,526 features $F$, in which each concept $w$ is assigned a feature vector $f(w)\in\mathbb{R}^{|F|}$: the value of $f(w)_{Q}$ is the value of the feature $Q$ for the word $w$ (see Table 4 in the Appendix).
@@
As the goal of this experiment was to evaluate the quality of their embeddings, the models were not fine-tuned for this task: following Rosenfeld and Erk (2022), we used the Modified Adsorption Talukdar and Crammer (2009) graph label-propagation algorithm to estimate properties, using the cosine distance between two words’ embeddings to weight the graph edge between them.
@@ 5.2 Results
The results of these experiments are given in Table 1. Although the two (original) BERT models outperform GFoLDS all four benchmarks, the two BERT comparison models both lag far behind our model, demonstrating that it is able to learn useful representations with less data than textual LMs across a wide range of downstream tasks.
@@ 6 Scalability
While the results of Section 5 show that GFoLDS is a viable competitor to textual models trained on similar amounts of data, this model is still outperformed by the original BERT models. It is therefore crucial to establish the scalability of GFoLDS: the degree to which we would expect its performance to scale if it were larger and/or pre-trained on more data. To that end, we applied the techniques of Muennighoff et al. (2024) to GFoLDS, to determine the degree to which GFoLDS is under- or over-parameterized—and therefore, by the scaling laws established in Muennighoff et al. (2024), over-/under-trained.
@@ 6.1 Experimental Setup
We pretrained five GFoLDS models on 50%, 25%, 12.5%, 6.25%, and 3.125% of the pretraining data used in Section 3.3. Following Muennighoff et al. (2024), we ensured that the iterations pretrained on less data always use a randomly-selected subset of the dataset used in those with more data. Aside from the differing number of pretraining tokens, all models were pretrained using the same procedure and hyperparameters as described in Section 3.3, as the focus of this experiment was the effect of training tokens with a fixed parameter count
@@ 6.2 Results
Figure 6 shows that final pretraining loss consistently decreases with the number of pretraining tokens from 3.125% to 50% of the data. After this point, the final loss value plateaus: the 50% run (1.3232) finishes with a lower cross-entropy loss than the actual (100%) run (1.3331), although this slight difference is likely explained by the noise introduced by random initialization of the models’ parameters. In Appendix H, we prove that—assuming that an analogue of the Muennighoff et al. (2024) scaling laws holds for GFoLDS—it can only be the case that the final loss for the 100% run is (roughly) equal to that for the 50% run if GFoLDS is underparameterized for both the 100% run and the 50% run. This is to say that GFoLDS requires much less pretraining data per parameter than textual models: the Chinchilla Scaling Laws Hoffmann et al. (2022) predict that a textual LLM with the same parameter count ($\sim$174 million) as GFoLDS necessitates $\sim$5 billion pretraining tokens—roughly twenty times more than the $\sim$254 million tokens at which GFoLDS becomes overparameterized.
@@ 7 Conclusion
To the best of our knowledge, the experiments in Section 5 represent the first time that a language model pretrained solely over logical forms has been evaluated on a wide range of downstream tasks. GFoLDS’ capacity to be applied to tasks ranging from RELPRON to SNLI (see Section 5) demonstrates this model’s versatility, and these results therefore represent a significant step towards demonstrating the practical viability of LFLMs.
@@ 8.1 Limitations of the Current Approach
Multiple-Sentence Inputs
@@ C.1 GFoLDS
During BERT’s pretraining procedure, while 15% of the input tokens are selected for prediction, only 80% of the selected tokens are masked: this is to account for the mismatch between the model’s pretraining and fine-tuning distributions that arises from the fact that the [MASK] token only occurs during pretraining. However, for GFoLDS, the [MASK] token does occur during fine-tuning as well, due to OOV items (as discussed in Section 3.2). Furthermore, Wettig et al. (2023) find that higher selection rates—and higher masking rates—result in improved performance on downstream tasks, when compared to the selection and masking/replacement rates reported in Devlin et al. (2019). For these reasons, we chose to mask 100% of the selected tokens during pre-training, with the slightly higher selection probability of 20%.
@@
We pretrained GFoLDS with a batch size of 16 for four epochs with the AdamW optimizer Loshchilov and Hutter (2017) and a weight decay value of $10^{-5}$. We set an initial learn rate of $10^{-5}$, with a linearly interpolated learn rate between values of $2\times 10^{-5}$ at the end of the first epoch, $10^{-5}$ at the end of the second, $3\times 10^{-6}$ at the end of the third, and $10^{-6}$ at the end of the fourth. That is to say that the learn rate increased linearly from $10^{-5}$ to $2\times 10^{-5}$ during the first epoch, decreased linearly from $2\times 10^{-5}$ to $10^{-5}$ during the second epoch, and so on.
@@ C.2 BERT Comparison Models
Appendix D RELPRON Task Setup
To evaluate the models on RELPRON, we constructed templates out of each (term, hypernym, property) triple: for example, the triple (telescope, device, astronomers use) yields the template “a device that astronomers use is a telescope” (see Table 2). We then replaced the target term with the [MASK] token (e.g. “a device that astronomers use is a [MASK]”): the probability assigned to a given term under the masked distribution is taken as proportional to the probability that the property applies to that term.
Appendix E SNLI Fine-Tuning Details
We fine-tuned GFoLDS and $\text{BERT}_{\text{base}}$ for five epochs with a batch size of 16, a weight decay value of $10^{-5}$, an initial learn rate of $10^{-5}$, and a linearly-interpolated learn rate (updated at each batch) between values of $2\times 10^{-5}$ at the end of the first epoch, $3\times 10^{-5}$ at the end of the third, $10^{-6}$ at the end of the fourth, and $10^{-7}$ at the end of the fifth. We fine-tuned $\text{BERT}_{\text{large}}$ with identical hyperparameters, except all learn rates mentioned above were multiplied by 1/10 for this model ($\text{BERT}_{\text{large}}$ was unstable with the higher learn rate used for GFoLDS and $\text{BERT}_{\text{base}}$).
@@ Appendix E SNLI Fine-Tuning Details
Appendix F MegaVeridicality V2.1 Task Setup
@@
Appendix G Elementary Tasks
@@
Appendix H Scalability
H.1 Background and Notation
@@ H.2 Proof
It is therefore reasonable to assume that $U_{D^{\prime}}\approx U_{D}/2$, where $U_{D}$ denotes the dataset for the 100% run, and $U_{D^{\prime}}$ that of the 50% run (see Section 6). We may express the relationship between the 50% run loss and the 100% loss in the notation introduced in Appendix H.1 (while leaving the exact definition of $U_{D}$ to future work): $L^{(4)}(N,U_{D})\approx L^{(4)}(N,U_{D}/2)$. Recall from Equation 5a that the term $E$ is a constant, and so can be ignored. We are then left with the following (approximate) equality in Equation 6a.

### v2

@@ 5 Downstream Tasks
In order to demonstrate the viability of LFLMs, we evaluated GFoLDS on a series of downstream tasks: our primary objective was to demonstrate that language models over logical forms are able to learn from less data than their textual counterparts.
To that end, we compared GFoLDS to the BERT comparison (BERT-C) models pretrained on the same data (see Section 4.1), in order to show that GFoLDS is able to outperform a textual LM when pretrained on the same amount of data: it then follows that GFoLDS would be able to perform as well as a textual LM with less pretraining data.
We additionally compared GFoLDS to the original $\text{BERT}_{\text{base}}$ and $\text{BERT}_{\text{large}}$ (uncased) models, both of which are pretrained on roughly 6.5 times more data than GFoLDS.
@@ 5.1 Datasets
We evaluated the models on four downstream tasks: the RELPRON task described in Section 4.2.1, SNLI Bowman et al. (2015), the MegaVeridicality V2.1 factuality dataset White et al. (2018), and the McRae et al. (2005) property inference dataset.
@@ 5.1.1 SNLI
All five models were fine-tuned end-to-end on SNLI with a two-layer, feed-forward MLP classifier over their mean-pooled token/node embeddings. The details of our SNLI fine-tuning hyperparameters are located in Appendix D.
Graph Representations.
@@
$\text{BERT}_{\text{base}}$ 0.667 0.690 90.7% 84.2% 0.247
$\text{BERT}_{\text{large}}$ 0.768 0.769 91.1% 85.6% 0.241
@@
[…]
@@
The objective in this task is to determine whether the event denoted by a subordinate clause is true, given the context of the matrix clause (see Table 3 in the Appendix). The MegaVeridicality V2.1 dataset consists of 5,026 examples, each with ten annotations of the factuality of the event: for this experiment, we converted the data into a binary classification task via majority voting of the annotated labels (see Appendix E).
@@
To evaluate the quality of GFoLDS’ embeddings in comparison to BERT, we evaluated the models on a property inference task using the McRae et al. (2005) feature norm database. This database consists of a set of 541 concepts (words) $W$ and 2,526 features $F$, where each concept $w\in W$ is assigned a feature vector $f(w)\in\mathbb{R}^{|F|}$: the value of $f(w)_{Q}$ is the value of the feature $Q$ for the word $w$ (see Table 4 in the Appendix).
@@
As the goal of this experiment was to evaluate the quality of their embeddings, the models were not fine-tuned for this task: again following Rosenfeld and Erk (2022), we used the Modified Adsorption Talukdar and Crammer (2009) graph label-propagation algorithm to estimate properties, using the cosine distance between two words’ embeddings to weight the graph edge between them.
@@ 5.2 Results
The results of these experiments are given in Table 1. Although the original BERT models outperform GFoLDS, the BERT comparison models trained on the same data as our model (BERT-C; see Section 4.1) both lag far behind GFoLDS on all four benchmarks, demonstrating across a wide range of downstream tasks that our model is able to learn useful representations with less data than its textual counterparts.
@@ 6 Scalability
While the results of Section 5 show that GFoLDS outperforms textual models trained on similar amounts of data, this model is still outperformed by the original BERT models. It is therefore crucial to establish the scalability of GFoLDS: the degree to which we would expect its downstream performance to scale if it were larger and/or pretrained on more data. To that end, we applied the techniques of Muennighoff et al. (2024) to GFoLDS, to determine the degree to which our model is under- or over-parameterized—and therefore, by the scaling laws established in Muennighoff et al. (2024), over- or under-trained.
@@ 6.1 Experimental Setup
We pretrained five GFoLDS models on 50%, 25%, 12.5%, 6.25%, and 3.125% of the pretraining data used in Section 3.3. Following Muennighoff et al. (2024), we ensured that the iterations pretrained on less data always use a randomly-selected subset of the dataset used in those with more data.
Aside from the differing number of pretraining tokens, all models were pretrained using the same procedure and hyperparameters as described in Section 3.3, as the focus of this experiment was the effect of pretraining tokens with a fixed parameter count.
@@ 6.2 Results
Figure 6 shows that final pretraining loss consistently decreases as the number of pretraining tokens increases from 3.125% to 50% of the data. After this point, the final loss value plateaus: the 50% run (1.3232) finishes with a slightly lower cross-entropy loss than the actual (100%) run (1.3331), although this minor difference is likely explained by the noise introduced by random initialization of the models’ parameters.
In Appendix G, we prove that—assuming that an analogue of the Muennighoff et al. (2024) scaling laws holds for the GFoLDS architecture—it can only be the case that the final loss for the 100% run is (roughly) equal to that for the 50% run if GFoLDS is underparameterized for both the 100% run and the 50% run. This is to say that GFoLDS requires much less pretraining data per parameter than textual models: the Chinchilla Scaling Laws Hoffmann et al. (2022) predict that a textual LLM with the same parameter count ($\sim$174 million) as GFoLDS necessitates $\sim$5 billion pretraining tokens—roughly twenty times more than the $\sim$254 million tokens for which GFoLDS is overparameterized.
@@ 7 Conclusion
To the best of our knowledge, the experiments in Section 5 represent the first time that a language model pretrained solely over logical forms has been evaluated on a wide range of downstream tasks. GFoLDS’ capacity to be applied to tasks ranging from RELPRON to SNLI—and consistently outperform BERT models pretrained on the same amount of data—demonstrates this model’s versatility. These results therefore represent a significant step towards demonstrating the practical viability of LFLMs.
@@ 8.1 Limitations of the Current Approach
Multiple-Sentence Inputs.
@@ 8.2 Experimental Limitations
Acknowledgments
Computational resources for these experiments were provided by the Center for Computational Research at the University at Buffalo .
@@ C.1 GFoLDS
During BERT’s pretraining procedure, while 15% of the input tokens are selected for prediction, only 80% of the selected tokens are masked: this is to account for the mismatch between the model’s pretraining and fine-tuning distributions that arises from the fact that the [MASK] token only occurs during pretraining. However, for GFoLDS, the [MASK] token does occur during fine-tuning as well, due to OOV items (as discussed in Section 3.2). Furthermore, Wettig et al. (2023) find that higher selection rates—and higher masking rates—result in improved performance on downstream tasks, when compared to the selection and masking/replacement rates reported in Devlin et al. (2019). For these reasons, we chose to mask 100% of the selected tokens during pretraining, with the slightly higher selection probability of 20%.
@@
Figure 10: GFoLDS pretraining learn rate schedule.
We pretrained GFoLDS with a batch size of 16 for four epochs with the AdamW optimizer Loshchilov and Hutter (2017) and a weight decay value of $10^{-5}$. We set an initial learn rate of $10^{-5}$, with a linearly interpolated learn rate between values of $2\times 10^{-5}$ at the end of the first epoch, $10^{-5}$ at the end of the second, $3\times 10^{-6}$ at the end of the third, and $10^{-6}$ at the end of the fourth (see Figure 10). That is to say that the learn rate increased linearly from $10^{-5}$ to $2\times 10^{-5}$ during the first epoch, decreased linearly from $2\times 10^{-5}$ to $10^{-5}$ during the second epoch, and so on.
@@ C.2 BERT Comparison Models
Appendix D SNLI Fine-Tuning Details
We fine-tuned GFoLDS and $\text{BERT}_{\text{base}}$ for five epochs with a batch size of 16, a weight decay value of $10^{-5}$, an initial learn rate of $10^{-5}$, and a linearly-interpolated learn rate (updated at each batch) between values of $2\times 10^{-5}$ at the end of the first epoch, $3\times 10^{-5}$ at the end of the third, $10^{-6}$ at the end of the fourth, and $10^{-7}$ at the end of the fifth (see Figure 15). We fine-tuned $\text{BERT}_{\text{large}}$ with identical hyperparameters, except all learn rates mentioned above were multiplied by 1/10 for this model ($\text{BERT}_{\text{large}}$ was unstable with the higher learn rate used for GFoLDS and $\text{BERT}_{\text{base}}$).
Figure 15: SNLI fine-tuning learn rate schedule for GFoLDS and the $\text{BERT}_{\text{base}}$ models.
@@ Appendix D SNLI Fine-Tuning Details
Appendix E MegaVeridicality V2.1 Task Setup
@@
Appendix F Elementary Tasks
@@
Appendix G Scalability
G.1 Background and Notation
@@ G.2 Proof
It is therefore reasonable to assume that $U_{D^{\prime}}\approx U_{D}/2$, where $U_{D}$ denotes the dataset for the 100% run, and $U_{D^{\prime}}$ that of the 50% run (see Section 6). We may express the relationship between the 50% run loss and the 100% loss in the notation introduced in Appendix G.1 (while leaving the exact definition of $U_{D}$ to future work): $L^{(4)}(N,U_{D})\approx L^{(4)}(N,U_{D}/2)$. Recall from Equation 5a that the term $E$ is a constant, and so can be ignored. We are then left with the following (approximate) equality in Equation 6a.